# VL-Nav: Real-time Vision-Language Navigation with Spatial Reasoning

### v1

@@ Abstract
Vision-language navigation in unknown environments is crucial for mobile robots. In scenarios such as household assistance and rescue, mobile robots need to understand a human command, such as “find a person wearing black”. We present a novel vision-language navigation (VL-Nav) system that integrates efficient spatial reasoning on low-power robots. Unlike prior methods that rely on a single image-level feature similarity to guide a robot, we introduce the heuristic-vision-language (HVL) spatial reasoning for goal point selection. It combines pixel-wise vision-language features and heuristic exploration to enable efficient navigation to human-instructed instances in various environments robustly. We deploy VL-Nav on a four-wheel mobile robot and conduct comprehensive navigation tasks in various environments of different scales and semantic complexities, indoors and outdoors. Remarkably, VL-Nav operates at a real-time frequency of 30 Hz with a Jetson Orin NX, highlighting its ability to conduct efficient vision-language navigation. Experimental results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and heuristic cues for mobile robots. (a) Hallway: The wheeled robot is tasked with “find a man in gray” in a hallway. Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the nearby frontier point as the goal and successfully locating the missing person. The value map shows that the “gray cloth” VL cue prioritizes the right-side area, marked by yellow square points. (b) Apartment: The robot is tasked with “Go to the tall white trash bin.” It detected two different-sized white trash bins in bottom camera observation. However, it assigns a higher confidence score (0.98) to the taller bin than the shorter one (0.48). These pixel-wise VL features are incorporated into the spatial distribution to select the correct goal point, guiding the robot toward the taller bin.
@@ I Introduction
To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer. VL-Nav employs heuristic vision-language (HVL) spatial reasoning on both frontier-based and instance-based target points. VL-Nav initially generates frontier-based target points from the dynamic occupancy map using partial frontier detection. It confines the search to a manageable field of view, thereby reducing computational overhead. Additionally, instance-based target points are incorporated to emulate human search patterns, enabling the robot to approach and verify potential target objects, thereby enhancing the success rate. To select the most informative goal point, we introduce HVL spatial reasoning. This technique first transforms pixel-wise vision-language features into a spatial scoring distribution using a Gaussian mixture model. Each target point is then assigned a vision-language semantic score based on this distribution. Subsequently, a curiosity-driven heuristic weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into HVL scores. The target point with the highest HVL score is ultimately selected as the goal point. This HVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
Once a goal is selected, VL-Nav utilizes a classic planner for real-time obstacle-aware path planning, enabling seamless adaptation to partially known environments. By combining pixel-wise vision-language features with heuristics via the novel HVL spatial reasoning, VL-Nav surpasses all the baselines, achieving human-like intelligent navigation while remaining computationally feasible for real-world deployment.
@@
  • We introduce the heuristic-vision-language (HVL) spatial reasoning combining pixel-wise vision-language features and heuristic exploration for efficient and robust VLN..
@@ II Related Works
Figure 2: An overview of the VL-Nav pipeline. VL-Nav processes inputs including prompts, RGB images, odometry poses, and LiDAR scans. The Vision-Language (VL) module conducts open-vocabulary pixel-wise detection to identify areas and objects related to the prompt, generating instance-based target points. Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map. Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool. VL-Nav employs heuristic-vision-language (HVL) spatial reasoning to select the most effective goal point from this pool for path planning.
@@
VL-Nav significantly advances vision-language navigation (VLN) by addressing the limitations of classical, end-to-end, and modular learning methods. It employs a rolling occupancy grid and partial frontier detection, reducing computational overhead and enhancing adaptability in dynamic environments on low-power platforms. The integration of instance-based target points with heuristic vision-language (HVL) spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection. Unlike systems reliant on heavy computational models, VL-Nav optimizes for resource-constrained platforms and maintains real-time performance at 30 Hz. This makes VL-Nav not only practical for real-world applications but also superior in navigating complex environments efficiently.
@@ III Methodology
This section presents our VL-Nav pipeline (Fig. 2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and Heuristic-Vision-Language (HVL) spatial reasoning to select and publish a navigation goal. This goal is not only aligned with the given prompt but also leads the robot to explore unknown areas more effectively.
@@ III-C Instance-Based Target Points (IBTP)
III-D Heuristic-Vision-Language (HVL) Spatial Reasoning
Once frontier centroids and instance-based targets are gathered, the system computes an HVL score for each candidate goal $\mathbf{g}$. Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance. The angular offset is
@@ III-D Heuristic-Vision-Language (HVL) Spatial Reasoning
Heuristic Cues:
We add two heuristic terms to guide navigation toward larger unexplored areas and prevent the system from repeatedly selecting faraway goals that offer only a marginal increase in VL score—thus reducing unnecessary back-and-forth movement which is extremely important for the large-scale environments like the hallway and the outdoor environments in our experiments.
@@
$S_{\mathrm{HVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights. Note that the VL score is assigned when a point is initially detected, whereas the heuristic terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated. For instance-based goals, these heuristic cues are omitted because their primary purpose is verification rather than exploration.
@@ III-E Select & Publish Goal Point
After evaluating HVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate. Algorithm 2 sketches the selection process, prioritizing instances first, then frontiers.
@@
0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $HVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
@@ III-F Path Planning with FAR Planner
Once an HVL goal is selected, we hand off point-goal path planning to the FAR Planner [34], which represents obstacles polygonally and updates a visibility graph in real time. This enables efficient re-planning in partially unknown environments, often outperforming search- or sampling-based methods. A local planner refines FAR Planner’s waypoints into short-horizon velocity commands, ensuring swift reactions to new obstacles and seamless integration with VL-Nav’s zero-shot vision-language objectives.
@@ IV-A Experimental Setting
We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [36], (3) VLNav without instance-based target points, (4) VLNav without heuristic terms, and (5) the full VLNav configuration. Because the original VLFM relies on BLIP-2 [14], which is too computationally heavy for real-time edge deployment, we use the YOLO-World [6] model instead to generate per-observation similarity scores for VLFM. Each method is tested under the same conditions to ensure a fair comparison of performance.
@@ IV-B Main Results
As reported in Table I, our full VL-Nav consistently obtains the highest Success Rate (SR) and Success weighted by Path Length (SPL) across all four environments. In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating HVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
@@
Heuristics Contributions:
The Heuristic Score is also significant to VL-Nav’s performance. It merges two key components:
@@
  • Unknown-Area Heuristic: Rewards navigation toward regions that yield more information.
Our ablations reveal that removing the distance-scoring element (VL-Nav w/o heuristics) degrades both SR and SPL, particularly in the more cluttered environments. Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of HVL addresses a complementary aspect of semantic navigation.
@@
Although the VLFM approach [36] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and HVL-based spatial reasoning. Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the heuristic cues via the HVL spatial reasoning. These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our HVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
@@
In conclusion, the experimental results confirm that VL-Nav delivers superior vision-language navigation across diverse, unseen real-world environments. By fusing frontier-based target points detection, instance-based target points, and the HVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency. The system’s robust performance, even in large or cluttered domains, highlights its potential as a practical solution for zero-shot vision-language navigation on low-power robots.
@@ V Discussion
By contrast, VL-Nav remains robust across all sizes. Although there is a slight decrease when moving from Mid (91.7% SR, 0.852 SPL) to Big (82.3% SR, 0.677 SPL), its overall performance remains substantially higher than the baselines. A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o Heuristic), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
These findings underscore two key factors. First, HVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig. 1) while also prioritizing areas with substantial unknown space, thus avoiding the extreme inefficiencies of frontier-based exploration faces in large, open environments. Second, instance-based target points (IBTP) enable more targeted navigation, sustaining high success rates even as map size grows.
@@
These results underscore that VL-Nav gains more pronounced benefits as the environment offers more semantic context. Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and heuristic exploration. The fusion of the HVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
@@ VI-B Conclusion
In this paper, we introduced VL-Nav, a vision-language navigation framework that operates efficiently in real time on resource-constrained platforms. By combining pixel-wise vision-language features with a heuristic exploration strategy, our HVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings. Empirically, VL-Nav not only achieves real-time navigation at 30 Hz on a Jetson Orin NX but also surpasses existing approaches by 44.15%, achieving an overall success rate of 86.3%. The key to these gains lies in effectively leveraging semantic cues from vision-language embeddings to prioritize frontiers and potential object instances, thereby aligning navigation decisions more closely with human-like reasoning.

### v2

@@ Abstract
Vision-language navigation in unknown environments is crucial for mobile robots. In scenarios such as household assistance and rescue, mobile robots need to understand a human command, such as “find a person wearing black”. We present a novel vision-language navigation (VL-Nav) system that integrates efficient spatial reasoning on low-power robots. Unlike prior methods that rely on a single image-level feature similarity to guide a robot, our method integrates pixel-wise vision-language features with curiosity-driven exploration. This approach enables robust navigation to human-instructed instances across diverse environments. We deploy VL-Nav on a four-wheel mobile robot and evaluate its performance through comprehensive navigation tasks in both indoor and outdoor environments, spanning different scales and semantic complexities. Remarkably, VL-Nav operates at a real-time frequency of 30 Hz with a Jetson Orin NX, highlighting its ability to conduct efficient vision-language navigation. Results show that VL-Nav achieves an overall success rate of 86.3%, outperforming previous methods by 44.15%.
Figure 1: We propose VL-Nav, a real-time zero-shot vision-language navigation approach with spatial reasoning that integrates pixel-wise vision-language features and curiosity-based exploration for mobile robots. (a) Hallway: The wheeled robot is tasked with “find a man in gray” in a hallway. Unlike the classical frontier-based method (red line) and VLFM (green line), VL-Nav (blue line) leverages pixel-wise vision-language (VL) features from the “gray cloth” cue for spatial reasoning, selecting the most VL-correlated goal point and successfully locating the missing person. The value map shows that the “gray cloth” VL cue prioritizes the right-side area, marked by yellow square points. (b) Apartment: The robot is tasked with “Go to the tall white trash bin.” It detected two different-sized white trash bins in bottom camera observation. However, it assigns a higher confidence score (0.98) to the taller bin than the shorter one (0.48). These pixel-wise VL features are incorporated into the spatial distribution to select the correct goal point, guiding the robot toward the taller bin.
@@ I Introduction
To bridge this gap, we present vision-language navigation (VL-Nav), a novel navigation framework optimized for low-power robots, achieving zero-shot VLN at 30 Hz with an onboard computer. VL-Nav employs the spatial reasoning on both frontier-based and instance-based target points. VL-Nav initially generates frontier-based target points from the dynamic occupancy map using partial frontier detection. It confines the search to a manageable field of view, thereby reducing computational overhead. Additionally, instance-based target points are incorporated to emulate human search patterns, enabling the robot to approach and verify potential target objects, thereby enhancing the success rate. To select the most informative goal point, we introduce CVL spatial reasoning. This technique first transforms pixel-wise vision-language features into a spatial scoring distribution using a Gaussian mixture model. Each target point is then assigned a vision-language semantic score based on this distribution. Subsequently, a curiosity-driven weighting, which encourages the robot to explore unknown areas, is applied to adjust these scores into CVL scores. The target point with the highest CVL score is ultimately selected as the goal point. This CVL spatial reasoning process ensures that the selected goal point not only closely aligns with the human description but also guides the robot to explore unknown areas.
Once a goal is selected, VL-Nav utilizes a classic planner for real-time obstacle-aware path planning, enabling seamless adaptation to partially known environments. By combining pixel-wise vision-language features with curiosity-based exploration via the novel CVL spatial reasoning, VL-Nav surpasses all the baselines, achieving intelligent navigation while remaining computationally feasible for real-world deployment.
@@
  • We empower VL-Nav with spatial reasoning by integrating pixel-wise vision-language features and curiosity-driven exploration for more efficient VLN..
@@ II Related Works
Figure 2: An overview of the VL-Nav pipeline. VL-Nav processes inputs including prompts, RGB images, odometry poses, and LiDAR scans. The Vision-Language (VL) module conducts open-vocabulary pixel-wise detection to identify areas and objects related to the prompt, generating instance-based target points. Concurrently, the map module performs terrain analysis and manages a dynamic occupancy map. Frontier-based target points are then identified based on this occupancy map, along with the instance points, forming a candidate points pool. VL-Nav employs spatial reasoning to select the most effective goal point from this pool for path planning.
@@
VL-Nav significantly advances vision-language navigation (VLN) by addressing the limitations of classical, end-to-end, and modular learning methods. It employs a rolling occupancy grid and partial frontier detection, reducing computational overhead and enhancing adaptability in dynamic environments on low-power platforms. The integration of instance-based target points with spatial reasoning enables pixel-wise vision-language feature interpretation and context-aware goal selection. Unlike systems reliant on heavy computational models, VL-Nav optimizes for resource-constrained platforms and maintains real-time performance at 30 Hz. This makes VL-Nav not only practical for real-world applications but also superior in navigating complex and various environments efficiently.
@@ III Methodology
This section presents our VL-Nav pipeline (Fig. 2), which leverages a rolling occupancy grid map, frontier-based and instance-based target points, and curiosity-vision-language (CVL) Scoring Policy to select and publish a navigation goal. This goal is not only aligned with the given prompt but also leads the robot to explore unknown areas more effectively.
@@ III-C Instance-Based Target Points (IBTP)
III-D CVL Scoring Policy
Once frontier centroids and instance-based targets are gathered, the system computes an CVL score for each candidate goal $\mathbf{g}$. Let $(x_{r},y_{r})$ be the robot position, $\theta_{r}$ the heading, and $d(\mathbf{x}_{r},\mathbf{g})$ the Euclidean distance. The angular offset is
@@ III-D CVL Scoring Policy
Curiosity Cues:
We add two curiosity terms to guide navigation toward larger unexplored areas and prevent the system from repeatedly selecting faraway goals that offer only a marginal increase in VL score—thus reducing unnecessary back-and-forth movement which is extremely important for the large-scale environments like the hallway and the outdoor environments in our experiments.
@@
$S_{\mathrm{CVL}}(\mathbf{g})\;=\;\Bigl{(}w_{\mathrm{dist}}\;S_{\mathrm{dist}}(\mathbf{g})\;+\;w_{\mathrm{VL}}\;S_{\mathrm{VL}}(\mathbf{g})\;\cdot\;S_{\mathrm{unknown}}(\mathbf{g})\Bigr{)},$
where $w_{\mathrm{VL}}$ and $w_{\mathrm{dist}}$ are scalar weights. Note that the VL score is assigned when a point is initially detected, whereas the curiosity terms (distance and unknown-area weighting) are evaluated at goal selection time—when the current distance and unknown ratio can be calculated. For instance-based goals, these curiosity cues are omitted because their primary purpose is verification rather than exploration.
@@ III-E Select & Publish Goal Point
After evaluating CVL scores for all instance-based and frontier goals, the system picks the highest-scoring candidate. Algorithm 2 sketches the selection process, prioritizing instances first, then frontiers.
@@
0:  Instance points $\{\mathbf{o}_{1},\dots\}$ with $VL\_score$, frontier points $\{\mathbf{f}_{1},\dots\}$ with $CVL\_score$, robot position $\mathbf{r}$, and threshold $\delta_{\mathrm{reached}}$ for discarding points too close to $\mathbf{r}$.
@@ III-F Path Planning with FAR Planner
Once an CVL goal is selected, we hand off point-goal path planning to the FAR Planner [34], which represents obstacles polygonally and updates a visibility graph in real time. This enables efficient re-planning in partially unknown environments, often outperforming search- or sampling-based methods. A local planner refines FAR Planner’s waypoints into short-horizon velocity commands, ensuring swift reactions to new obstacles and seamless integration with VL-Nav’s zero-shot vision-language objectives.
@@ IV-A Experimental Setting
We evaluate our approach in real-robot experiments against five methods: (1) classical frontier-based exploration, (2) VLFM [36], (3) VLNav without instance-based target points, (4) VLNav without curiosity terms, and (5) the full VLNav configuration. Because the original VLFM relies on BLIP-2 [14], which is too computationally heavy for real-time edge deployment, we use the YOLO-World [6] model instead to generate per-observation similarity scores for VLFM. Each method is tested under the same conditions to ensure a fair comparison of performance.
@@ IV-B Main Results
As reported in Table I, our full VL-Nav consistently obtains the highest Success Rate (SR) and Success weighted by Path Length (SPL) across all four environments. In particular, VL-Nav outperforms classical exploration by a large margin, confirming the advantage of integrating CVL spatial reasoning with partial frontier-based search rather than relying solely on geometric exploration.
@@
Curiosity Contributions:
The curiosity Score is also significant to VL-Nav’s performance. It merges two key components:
@@
  • Unknown-Area Weighting: Rewards navigation toward regions that yield more information.
Our ablations reveal that removing the distance-scoring element (VL-Nav w/o curiosity) degrades both SR and SPL, particularly in the more cluttered environments. Meanwhile, dropping the instance-based target points (IBTP) similarly lowers performance, reflecting how each piece of CVL addresses a complementary aspect of semantic navigation.
@@
Although the VLFM approach [36] harnesses vision-language similarity value, it lacks the pixel-wise vision-language features, instance-based target points verification mechanism, and CVL-based spatial reasoning. Consequently, VL-Nav surpasses VLFM in both SR and SPL by effectively combining the pixel-wise vision language features and the curiosity cues via the CVL spatial reasoning. These gains are especially pronounced in semantic complex (Apartment) and open-area (Outdoor) environments, underscoring how our CVL spatial reasoning enhance vision-language navigation in complex settings and scenarios.
@@
In conclusion, the experimental results confirm that VL-Nav delivers superior vision-language navigation across diverse, unseen real-world environments. By fusing frontier-based target points detection, instance-based target points, and the CVL spatial reasoning for goal selection, VL-Nav balances semantic awareness and exploration efficiency. The system’s robust performance, even in large or cluttered domains, highlights its potential as a practical solution for zero-shot vision-language navigation on low-power robots.
@@ V Discussion
By contrast, VL-Nav remains robust across all sizes. Although there is a slight decrease when moving from Mid (91.7% SR, 0.812 SPL) to Big (82.3% SR, 0.655 SPL), its overall performance remains substantially higher than the baselines. A similar trend appears for the ablated VL-Nav variants (w/o IBTP, w/o curiosity), which still surpass Frontier Exploration but fall behind the full VL-Nav system.
These findings underscore two key factors. First, CVL spatial reasoning allows our method to leverage fine-grained vision-language cues (e.g., the “gray cloth” in Fig. 1) while also prioritizing areas with substantial unknown space, thus avoiding the extreme inefficiencies of frontier-based exploration faces in large, open environments. Second, instance-based target points (IBTP) enable more targeted navigation, sustaining high success rates even as map size grows.
@@
These results underscore that VL-Nav gains more pronounced benefits as the environment offers more semantic context. Rather than being overwhelmed by clutter or complexity, the system capitalizes on language-guided instance-based detections and curiosity-based exploration . The fusion of the CVL spatial reasoning and instance-based target points provides a scalable strategy to locate human-specified targets, even in highly complex indoor domains.
@@ VI-B Conclusion
In this paper, we introduced VL-Nav, a vision-language navigation framework that operates efficiently in real time on resource-constrained platforms. By combining pixel-wise vision-language features with a curiosity-based exploration strategy, our CVL spatial reasoning method demonstrates robust performance in a variety of indoor and outdoor settings. Empirically, VL-Nav not only achieves real-time navigation at 30 Hz on a Jetson Orin NX but also surpasses existing approaches by 44.15%, achieving an overall success rate of 86.3%. The key to these gains lies in effectively leveraging semantic cues from vision-language embeddings to prioritize frontiers and potential object instances, thereby aligning navigation decisions more closely with human-like reasoning.